# Critical Dynamics of Holographic Superfluids Aristomenis Donos and Polydoros Kailidis

### v1

@@ 1 Introduction
Superfluid states of matter were first realised holographically in [15] and [16]. In the gravitational picture, the thermal state is modeled by a black hole with Hawking temperature $T$, corresponding to the field theory temperature. Below a certain $T=T_{c}$, a second branch of black hole solutions with scalar hair develops. Assuming that this new family of solutions has a lower free energy, a phase transition takes place. Since their introduction, holographic superfluids have been studied extensively. Most of the investigations of their hydrodynamic regime have been numerical in nature (see, e.g., [17, 18]), with the exception of some notable analytic treatments [19].
@@ 3 Thermodynamic fluctuations near criticality
The overall scale in the above two perturbations is fixed so that $\delta_{T}g_{tt}=-4\pi\,r+\mathcal{O}(r^{2})$, near the horizon, and $\delta_{\mu}B_{t}=1+\mathcal{O}(\frac{1}{r+R})$, near the boundary. At this point, let us introduce a bit of notation for later use. If $\delta_{st.}\in\{\delta_{T},\,\delta_{\mu},\,\delta_{\rho_{s}},\,\delta_{v_{i}},\,\delta_{m_{i}},\,\delta_{s_{ab}}\}$ and $f$ is one of the radial functions parameterising the bulk fields, then $\delta_{st.}f_{[2n]}$ denotes the $n+1$-thth term in the $\varepsilon$ expansion of $\delta_{st.}f$, including factors of $\varepsilon$.
@@ 4.1 Spacetime-dependent perturbations
due to the translational symmetry of the background spacetime (2). This reduces the problem to a system of linear ODEs for the radial functions $\delta f(r)$. By choosing the fixed function $S(r)$ to approach $S(r)\approx\frac{\ln r}{4\pi T}+\cdots$ close to the black hole horizon at $r=0$, we are guaranteed to have the correct infalling boundary conditions, provided that $\delta f(r)$ remains analytic there. Finally, in order for the boundary theory information to be packaged entirely in the asymptotics of $\delta f(r)$, we will choose $S(r)$ to behave as $S(r)=\mathcal{O}\left({(r+R)^{-4}}\right)$ close to the conformal boundary.
@@ 4.2 Building the hydrodynamic perturbation
In this subsection, we will focus on the hydrodynamic limit of the bulk theory. To achieve this, we will scale the frequency and wave-vector of the perturbation (4.1) according to $w=\omega\,\varepsilon^{2},\,q_{i}=k_{i}\,\varepsilon^{2}$, with $\omega,\,k_{i}$ of order $\mathcal{O}(\varepsilon^{0})$. Similarly to the expansion of the background and of its static fluctuations in powers of $\varepsilon$, we can assume that the radial function $\delta f(r)$ in equation (4.1) is expandable in $\varepsilon$ as well. In particular, we expand the amplitude fluctuation as,
@@
and all the rest of the bulk fields as [^10] The fact that $\delta g_{rt},\,\delta g_{ri},\,\delta B_{r}$ are zero at order $\mathcal{O}(\varepsilon^{0})$ is a (partial) gauge fix of the bulk diffeomorphisms.,
@@
By inspecting the expanded equations of motion in $\varepsilon$, we can show that the functions $\delta\rho_{[1]}(r)$ and $\delta f_{[2]}(r)$ must solve the same set of ODEs with the leading parts of the thermodynamic fluctuations we discussed in Section 3. Hence, we can decompose a generic bulk field $f$ in the corresponding linear basis of spacetime-independent fluctuations according to,
$\delta f_{[2]}\,\varepsilon^{2}=\delta_{T}f_{[0]}\,\delta T+\delta_{\mu}f_{[0]}\,\delta\mu+\delta_{v_{i}}f_{[0]}\,\delta v_{i}+\delta_{\rho_{s}}f_{[0]}\,\delta\pi+\delta_{s_{ab}}f_{[0]}\,\delta s_{ab}+\delta_{m_{i}}f_{[0]}\,\delta m_{i}\,$, (4.11)
with a similar expression for $\delta\rho_{[1]}\,\varepsilon$. Note that all of the coefficients $\delta T$, $\delta\mu$, $\delta v_{i}$, $\delta s_{ab}$, $\delta s_{a}$ and $\delta m_{i}$ are of order $\mathcal{O}(\varepsilon^{2})$, while $\delta\pi$ is of order $\mathcal{O}(\varepsilon^{3})$. Moreover, we can expand the gauge invariant combination $\delta m_{a}$ according to $\delta m_{a}=\delta m_{a[2]}\varepsilon^{2}+\delta m_{a[4]}\varepsilon^{4}+\mathcal{O}(\varepsilon^{6})$.
The coefficients $\delta T,\,\delta\mu,\,\delta v_{i}$ are naturally interpreted as the temperature, chemical potential and normal fluid velocity fluctuations of the dual field theory. On the other hand, the coefficient $\delta\pi$ multiplying $\delta_{\rho_{s}}f_{[0]}$, is not an external source for the amplitude. In particular, we note that $\delta_{\rho_{s}}\rho_{[0]}=\frac{\partial_{\rho_{s}}\rho_{[0]}}{\varepsilon^{2}}$, which is a multiple of the critical mode (see (3.19)) and hence is sourceless. The time-dependent amplitude source arises from the asymptotic behaviour of the subleading part of the amplitude, $\delta\rho_{[3]}=\frac{\delta\rho_{[3](s)}}{(r+R)^{3-\Delta\psi}}+\cdots$ instead; however, the external source $\delta\rho_{[3](s)}$ and the thermodynamic parameter $\delta\pi$ appearing in (4.11) are completely independent [^11] The parameter $\delta\pi$ is essentially the holographic counterpart of $\delta\pi$ in [25], which justifies our notation..
The expectation value for the amplitude is,
@@ 6.2 Quasinormal modes
We have numerically constructed these modes in the bulk for several values of the wavevector $q$, at fixed temperature $\frac{T}{T_{c}}\approx 0.9999922$, corresponding to $\varepsilon^{2}\approx\frac{T-T_{c}}{T_{c}}\approx 10^{-5}$. Figure 1 shows $\frac{\partial}{\partial q}\mathrm{Re}[\omega_{FS,+}]$ and $\frac{1}{2}\frac{\partial^{2}}{\partial q^{2}}\mathrm{Im}[\omega_{FS,+}]$ for the first sound mode. As is evident from the plots, this mode interpolates between the following asymptotic behaviours: For small $\frac{q}{\mu}$, it is the first sound of the superfluid phase, with $\omega\simeq 0.70404\,q-i\,45.5\,q^{2}$ and for large $\frac{q}{\mu}$ it behaves like the first sound of the normal phase, with $\omega\simeq 0.70525\,q-i\,0.1602\,q^{2}$. Notice that for small momenta the coefficient of the $q^{2}$ term, i.e. the attenuation constant, is significantly large (compared, e.g. to the corresponding speed of sound), in accordance with our remark in Section 3.1 of [25] that it diverges close to the critical point.

### v3

@@ 1 Introduction
Superfluid states of matter were first realised holographically in [15] and [16]. In the gravitational picture, the thermal state is modeled by a black hole with Hawking temperature $T$, corresponding to the field theory temperature. Below a certain $T=T_{c}$, a second branch of black hole solutions with scalar hair develops. Assuming that this new family of solutions has a lower free energy, a phase transition takes place. Since their introduction, holographic superfluids have been studied extensively. Most of the investigations of their hydrodynamic regime have been numerical in nature (see, e.g., [17, 18, 19]), with the exception of some notable analytic treatments [20].
@@ 3 Thermodynamic fluctuations near criticality
Following the discussion in [22] (see also Subsection 4.1 in this paper) it is convenient to introduce a function $S(r)$ behaving as $S(r)\approx\frac{\mathrm{ln}r}{4\pi T}+\cdots$, close to the horizon. In this way, the combination $t+S(r)$ becomes regular on the black hole horizon. We will also choose the asymptotics of $S(r)$ to approach $\mathcal{O}(1/r^{4})$ close to the boundary, convenient for the holographic dictionary. It is also worth noting that $S$ follows our expansion scheme, $S=S_{[0]}+\mathcal{O}(\varepsilon^{2})$, since it depends on the temperature.
@@
The overall scale in the above two perturbations is fixed so that $\delta_{T}g_{tt}=-4\pi\,r+\mathcal{O}(r^{2})$, near the horizon, and $\delta_{\mu}B_{t}=1+\mathcal{O}(\frac{1}{r+R})$, near the boundary.
At this point, let us briefly comment on the notation. In the above, we have expanded the derivatives of the background fields with respect to $T$ and $\mu$ along the curve (3). For instance, $\partial_{T}U=\partial_{T}U_{[0]}+\partial_{T}U_{[2]}\,\varepsilon^{2}+\cdots,$ and $\partial_{T}\rho=\frac{1}{\varepsilon}\left(\partial_{T}\rho_{[0]}+\partial_{T}\rho_{[2]}\,\varepsilon^{2}+\cdots\right)$. Notice that the factors $\partial_{T}U_{[2n]},\partial_{T}\rho_{[2n]},etc$ are order $\varepsilon^{0}$. Comparing with the expansions (3) we can identify e.g. $\delta U_{\ast(2)}=\partial_{T}U_{[0]}\delta T_{\ast(2)}+\partial_{\mu}U_{[0]}\delta\mu_{\ast(2)}+\partial_{\phi_{s}}U_{[0]}\delta\phi_{s\ast(2)}$. In addition, in the following, if $\delta_{st.}\in\{\delta_{T},\,\delta_{\mu},\,\delta_{\rho_{s}},\,\delta_{v_{i}},\,\delta_{m_{i}},\,\delta_{s_{ab}}\}$ and $f$ is one of the radial functions parameterising the bulk fields, then $\delta_{st.}f_{[2n]}$ will denote the $n+1$-thth term in the $\varepsilon$ expansion of $\delta_{st.}f$, including factors of $\varepsilon$.
@@ 4.1 Spacetime-dependent perturbations
due to the translational symmetry of the background spacetime (2). This reduces the problem to a system of linear ODEs for the radial functions $\delta f(r)$. The function $S(r)$ is the same as the one we introduced in the previous section. Its behaviour near the horizon guarantees the correct infalling boundary conditions (provided that $\delta f(r)$ remains analytic there), whereas its behaviour close to the conformal boundary ensures that the boundary theory information is packaged entirely in the asymptotics of $\delta f(r)$.
@@ 4.2 Building the hydrodynamic perturbation
In this subsection, we will focus on the hydrodynamic limit of the bulk theory. This means that we wish to capture the response of the system at small frequencies and long wavelengths compared to e.g. the temperature. However, in the nearly critical system that we consider, there is another small parameter at our disposal, namely the parameter $\varepsilon$ of the previous section, which quantifies the distance from the critical point. As explained in [26], to efficiently describe the coupled dynamics of the order parameter with the hydrodynamic degrees of freedom, we will need to connect all the small parameters available, through the relations $w=\omega\,\varepsilon^{2},\,q_{i}=k_{i}\,\varepsilon^{2}$, with $\omega,\,k_{i}$ of order $\mathcal{O}(\varepsilon^{0})$. Similarly to the expansion of the background and of its static fluctuations in powers of $\varepsilon$, we can assume that the radial function $\delta f(r)$ in equation (4.1) is expandable in $\varepsilon$ as well. In particular, we expand the amplitude fluctuation as,
@@
and all the rest of the bulk fields as [^10] The scaling of the hydrodynamic fluctuations with $\varepsilon$ is dictated by consistency with the equations of motion. Also, the fact that $\delta g_{rt},\,\delta g_{ri},\,\delta B_{r}$ are zero at order $\mathcal{O}(\varepsilon^{0})$ is a (partial) gauge fix of the bulk diffeomorphisms.,
@@
By inspecting the expanded equations of motion in $\varepsilon$, we can show that the functions $\delta\rho_{[1]}(r)$ and $\delta f_{[2]}(r)$ must solve the same set of ODEs with the leading parts of the thermodynamic fluctuations we discussed in Section 3. Hence, we can decompose them in the corresponding linear basis of spacetime-independent fluctuations according to,
$\delta f_{[2]}\,\varepsilon^{2}=\delta_{T}f_{[0]}\,\delta T+\delta_{\mu}f_{[0]}\,\delta\mu+\delta_{v_{i}}f_{[0]}\,\delta v_{i}+\delta_{\rho_{s}}f_{[0]}\,\delta\pi+\delta_{s_{ab}}f_{[0]}\,\delta s_{ab}+\delta_{m_{i}}f_{[0]}\,\delta m_{i}\,$,
$\delta\rho_{[1]}\,\varepsilon=\delta_{T}\rho_{[0]}\,\delta T+\delta_{\mu}\rho_{[0]}\,\delta\mu+\delta_{v_{i}}\rho_{[0]}\,\delta v_{i}+\delta_{\rho_{s}}\rho_{[0]}\,\delta\pi+\delta_{s_{ab}}\rho_{[0]}\,\delta s_{ab}+\delta_{m_{i}}\rho_{[0]}\,\delta m_{i}\,.$ (4.11)
Note that all of the coefficients $\delta T$, $\delta\mu$, $\delta v_{i}$, $\delta s_{ab}$, $\delta s_{a}$ and $\delta m_{i}$ are of order $\mathcal{O}(\varepsilon^{2})$, while $\delta\pi$ is of order $\mathcal{O}(\varepsilon^{3})$. Moreover, we can expand the gauge invariant combination $\delta m_{a}$ according to $\delta m_{a}=\delta m_{a[2]}\varepsilon^{2}+\delta m_{a[4]}\varepsilon^{4}+\mathcal{O}(\varepsilon^{6})$.
The coefficients $\delta T,\,\delta\mu,\,\delta v_{i}$ are naturally interpreted as the temperature, chemical potential and normal fluid velocity fluctuations of the dual field theory. On the other hand, the coefficient $\delta\pi$ is not an external source for the amplitude [^11] The parameter $\delta\pi$ is essentially the holographic counterpart of $\delta\pi$ in [26], which justifies our notation.: the corresponding term, $\delta_{\rho_{s}}\rho_{[0]}\,\delta\pi=\frac{\partial_{\rho_{s}}\rho_{[0]}}{\varepsilon^{2}}\,\delta\pi$, entering the expression (4.2) for $\delta\rho_{[1]}\,\varepsilon$, is a multiple of the critical mode (see (3.19)) and hence is sourceless.
It is important to stress that the subleading parts of the hydrodynamic fluctuations, $\delta\rho_{[3]}\varepsilon^{3}$ and $\delta f_{[4]}\varepsilon^{4}$, are not fixed solely by thermodynamics anymore, but are intimately related to the spacetime dependence of the problem. Specifically, they obey certain inhomogeneous ODEs, in which factors of $\omega$ and $k_{i}$ appear explicitly.
We choose to turn on an external source for $\delta\rho$ through the asymptotics of its subleading part, namely we take $\delta\rho_{[3]}=\frac{\delta\rho_{[3](s)}}{(r+R)}+\frac{\delta\rho_{[3](v)}}{(r+R)^{2}}+\cdots,$ and thus, $\delta\rho_{(s)}=\delta\rho_{[3](s)}\varepsilon^{3}\,.$ The expectation value for the amplitude is,
@@ 6.2 Quasinormal modes
We have numerically constructed these modes in the bulk for several values of the wavevector $q$, at fixed temperature [^19] The next-to-leading order effective theory we have found analytically in the previous sections is valid for a range of temperatures close to the critical point. We choose a temperature this close to $T_{c}$ to perform our numerics only because the interpolating behaviour of the quasinormal modes is more pronounced then, as can be clearly seen in the figures below. $\frac{T}{T_{c}}\approx 0.9999922$, corresponding to $\varepsilon^{2}\approx\frac{T-T_{c}}{T_{c}}\approx 10^{-5}$. Figure 1 shows $\frac{\partial}{\partial q}\mathrm{Re}[\omega_{FS,+}]$ and $\frac{1}{2}\frac{\partial^{2}}{\partial q^{2}}\mathrm{Im}[\omega_{FS,+}]$ for the first sound mode. As is evident from the plots, this mode interpolates between the following asymptotic behaviours: For small $\frac{q}{\mu}$, it is the first sound of the superfluid phase, with $\omega\simeq 0.70404\,q-i\,45.5\,q^{2}$ and for large $\frac{q}{\mu}$ it behaves like the first sound of the normal phase, with $\omega\simeq 0.70525\,q-i\,0.1602\,q^{2}$. Notice that for small momenta the coefficient of the $q^{2}$ term, i.e. the attenuation constant, is significantly large (compared, e.g. to the corresponding speed of sound), in accordance with our remark in Section 3.1 of [26] that it diverges close to the critical point.
@@
We would like now to focus on the gap of the Higgs mode within our effective theory. As shown in [26], the gap is given by, [^20] Where, $\tilde{\nu}_{\rho\rho}\equiv\left(\frac{\partial\rho_{v}}{\partial\rho_{s}}\right)_{s,\varrho}$.
$\omega_{g}=-i\frac{2\,\mathrm{Re}\Gamma_{0}+\mathrm{Re}Z_{n}\,\rho_{v}^{2}}{2\,\tilde{\nu}_{\rho\rho}+\mathrm{Re}Z_{\pi}}+\mathcal{O}(\varepsilon^{6})\,.$ (6.14)
We note that this formula captures both the leading ($\sim\varepsilon^{2}$) and the next-to-leading ($\sim\varepsilon^{4}$) behaviour of the gap. Hence, it improves and extends the validity of the corresponding formula found in [23] (and later in [24]),
$\omega_{g}=-i\frac{\,\mathrm{Re}\Gamma_{0}^{lead.}}{\,\tilde{\nu}_{\rho\rho}}+\mathcal{O}(\varepsilon^{4})\,$, (6.15)
with,
$\Gamma_{0}^{lead.}=\frac{\rho_{v}^{2}}{\varpi_{1}-i\varpi_{2}}\,$, (6.16)
which is correct only to leading order in $\varepsilon$.
In Figure 4 we plot the ratio $\frac{\omega_{g(Analytic)}}{\omega_{g(Numeric)}}$ for a range of temperatures close to the critical one. The Higgs gap $\omega_{g(Numeric)}$ was calculated numerically. For the blue line, the gap $\omega_{g(Analytic)}$ was calculated using the new formula (6.14), while for the red line, the gap $\omega_{g(Analytic)}$ was calculated using the old formula (6.15). We clearly see in the diagram that the new formula approximates the gap with much better accuracy than the old one, as we move away from $T_{c}$, as expected. Since the coefficient $Z_{\pi}$ features in the expression (6.14) for the gap, this numerical check validates the quantitative importance of this novel coefficient in the accurate description of the low energy spectrum.
Figure 4: Plot of $\frac{\omega_{g(Analytic)}}{\omega_{g(Numeric)}}$ for a range of temperatures close to the critical one. The gap $\omega_{g(Numeric)}$ was calculated numerically. The gap $\omega_{g(Analytic)}$ of the blue line was calculated with (6.14), and $\omega_{g(Analytic)}$ of the red line was calculated using (6.15).